OpenStudy (anonymous):

X2+Y2=100 when xy=18 what is the value of x and y

OpenStudy (anonymous):

\[\large (x+y)^2 = (x^2 + y^2) + 2(xy)\]

OpenStudy (anonymous):

Find x + y first..

OpenStudy (lgbasallote):

i wonder what water's plan is -_- guess i'll check it out

OpenStudy (anonymous):

x = -100, y = -1, Y = ±8 x = -10, y = -10, Y = ±8

OpenStudy (anonymous):

This is by wolfram alpha..

OpenStudy (lgbasallote):

wolfram doesnt really give right answers...if you dont know how to input it right...

OpenStudy (anonymous):

u can see here

OpenStudy (anonymous):

\[(x+y)^2 = 136 \implies (x+y)= \pm \sqrt{136}\]

OpenStudy (lgbasallote):

here comes a monologue solution >:))

OpenStudy (anonymous):

Im so lost right now -_-

OpenStudy (anonymous):

Now : \[x = \sqrt{136} - y\] \[(\sqrt{136} - y)y = 18\]

OpenStudy (lgbasallote):

i agree @Seamonkey0912 i think @waterineyes should go slow dont you agree? ;)

OpenStudy (anonymous):

And now I am also lost..

OpenStudy (anonymous):

Let me try first..

OpenStudy (lgbasallote):

maybe you should've done it the traditional way...

OpenStudy (anonymous):

\[y^2 - \sqrt{136}y + 18 = 0 \implies y = \frac{\sqrt{136} \pm 8}{2}\]

OpenStudy (lgbasallote):

x^2 + y^2 = 100 xy = 18 ^from there i can say x = 18/y now substitute this value of x into x^2 + y^2 =100 giving you (18/y)^2 + y^2 = 100 18^2/y^2 + y^2 = 100 324/y^2 + y^2 = 100 now multiply all terms by y^2 324 + y^4 = 100y^2 y^4 - 100y^2 + 324 = 0 use quadratic formula to get the value of y

OpenStudy (anonymous):

x^2+y^2 is not equal to (x+y)^2

OpenStudy (anonymous):

How do you know and where I have written this ??

OpenStudy (anonymous):

Ahhh o_O still lost I'm sorry

OpenStudy (anonymous):

I used: \[(x+y)^2 = x^2 + y^2 + 2xy \]

OpenStudy (anonymous):

Are you having answers with you @Seamonkey0912

OpenStudy (anonymous):

I thought it was +8, -8 but I can't get it to equal to 100

OpenStudy (anonymous):

I got: \[y = 4 \pm \sqrt{34}\]

OpenStudy (anonymous):

Now find x : \[x = 2 \sqrt{34} - 4 \pm \sqrt{34} \implies x = 3\sqrt{34} - 4 \quad Or \quad \sqrt{34} -4\]

OpenStudy (anonymous):

Similarly: \[x = -2 \sqrt{34} - y\] \[y^2 + 2\sqrt{34}y + 18 = 0 \implies y = -\sqrt{34} \pm 4\] And now the x again..

OpenStudy (phi):

using the idea \[ (x+y)^2 = x^2+2xy+y^2 \] and the given values \[ x^2+y^2= 100 \text{ and } 2xy=36 \] \[ (x+y)^2= 136 \] or \[ x+y= 2\sqrt{34} \] also \[ xy= 18 \] one way to solve is replace y= 18/x in the first equation to get \[ x+\frac{18}{x}-2\sqrt{34} =0 \] \[ x^2 -2\sqrt{34}x + 18= 0 \] use the quadratic formula to get \[ x= \sqrt{34} \pm 4 \] that means \[ x= \sqrt{34}+4 \text{ and } y= \sqrt{34}-4 \] check: \[ xy= ( \sqrt{34}+4)( \sqrt{34}-4)= 34-16= 18 \] and we will see that \[ x^2+y^2= 100 \]